### 14 Jan 2024 v2

# On the Conditions for a Quantum Violent Relaxation

Abstract
In general, classical fully-connected systems are known to undergo violent relaxation. This phenomenon refers to the relaxation of observables to stationary, non-thermal, values on a finite timescale, despite their long-time dynamics being dominated by mean-field effects in the thermodynamic limit. Here, we analyze the “quantum" violent relaxation by studying the dynamics of generic many-body systems with two-body, all-to-all, interactions in the thermodynamic limit. We show that, in order for violent relaxation to occur very specific conditions on the spectrum of the mean-field effective Hamiltonian have to be met. These conditions are hardly met and “quantum" violent relaxation is observed rarely with respect to its classical counterpart. Our predictions are validated by the study of a spin model which, depending on the value of the coupling, shows a transition between violent-relaxation and a generic prethermal phase. We also analyze a spin version of the quantum Hamiltonian-Mean-Field model, which is shown not to exhibit violent-relaxation. Finally, we discuss how the violent-relaxation picture emerges back in the classical limit. Our results demonstrate how, even in the mean-field regime, quantum effects have a rather dramatic impact on the dynamics, paving the way to a better understanding of light-matter coupled systems.
Introduction: The physics of long-range interacting systems [1, 2], i.e. those systems whose microscopic components are coupled through slow-decaying power-law potentials $\sim r^{-\alpha}$, has recently experienced a new wave of interest due to their possible realization in multiple experimental setups [3]. Depending on the decay exponent $\alpha$, long-range interactions generate several novel features in classical physics, such as tuneable critical scaling in the weak long-range regime ($\alpha>d$) [4, 5, 6, 7] or ensemble inequivalece [1, 8, 9] and quasi-stationary states [10] in the strong long-range regime ($\alpha<d$). Quantum Long-range physics has been shown to be even richer, displaying plenty of peculiar features such as anomalous defect scaling [11, 12, 13, 14], long-lived metastable states [15, 16, 17], time crystals [18, 19, 20], dynamical phase transitions [20, 21] and several others [3, 22].
Violent relaxation is a characteristic feature of classical long-range systems out of equilibrium [23, 1]. This phenomenon has been observed in fully-connected models, including the paradigmatic Hamiltonian-Mean-Field (HMF) model [24, 1], and it is believed to be a key ingredient to explain the formation of large-scale astrophysical structures [25, 26, 27, 28, 10, 29, 30, 31]. The HMF dynamics, in the thermodynamic limit, becomes mean-field, thus preventing the loss of memory and the onset of the thermalization. In spite of this, on a finite (size-independent) time-scale, the system is usually observed to relax toward a so-called quasi-stationary state, in which large-scale observables take a constant (non-thermal) value. While the general mechanism behind the phenomenon is generally understood [2, 26, 28, 32, 33, 31], no general framework, comparable to the H-theorem in the usual kinetic theory [34], has emerged yet.
In the quantum regime, fully connected Hamiltonians, akin to the HMF, can be engineered in cavity experiments [17, 35] or Rydberg gases [36, 37] or seen as a large-dimension approximation of short-range models [38, 39]. These models display a rich out-of-equilibrium phenomenology, including dynamical phase transitions [40, 41, 42, 43], anomalous entanglement scaling [44, 45, 46], Floquet time-crystals [47, 48, 49], which has proven to be robust against external noise and competing short-range perturbations [15, 50, 16, 51]. As the mean-field description becomes exact [52, 53, 54, 55], according to the general wisdom, the dynamics of fully-connected systems can be reduced to a small number of classical macroscopic degrees of freedom in the thermodynamic limit [41, 42], resulting in long-lived periodic oscillations of the order parameter (a picture easily generalized to most strong-long range systems [16]). This is the case, e.g., of the paradigmatic Lipkin-Meshkov-Glick (LMG) model [56, 43].
In spite of the fundamentally classical nature of the mean-field dynamics, the resulting phenomenology lies outside the violent-relaxation paradigm, proper of classical model in mean-field regime, suggesting that the non-thermal relaxation is hindered by quantum effects. Quantum versions of the HMF model have been introduced both in the fermionic [57] and bosonic [58] cases, and the possibility of violent-relaxation in the latter have been recently investigated in Ref. [59]. In spite of these preliminary results, a general picture to explain the onset of violent-relaxation in the quantum world is still missing. In this Letter we address the matter in its generality, finding the conditions under which fully-connected models undergo violent relaxation. We find that such a possibility is closely linked to the property of the spectrum of the mean-field effective Hamiltonian. This result forms a bridge between the classical and the quantum realms and highlights how, already at the level of mean-field dynamics, single-particle quantum effects do play an important role and can deeply affect the out-of-equilibrium phenomena.
Results: Let us consider a generic set of $N$ quantum spins coupled by a fully-connected, two-body, interaction which can be described by the Hamiltonian
$H=\sum_{j}H_{0}(\tau_{j})-\frac{1}{2N}\sum_{i,j}H_{1}(\tau_{i})H_{1}(\tau_{j})$ (1)
where $\tau_{j}$ denotes a set of quantum variable relative to the site $j$, such that $[\tau_{j},\tau_{l}]=0$ for $i\neq l$, and the $1/N$ factor is chosen in agreement with the so-called Kac scaling, which ensures an extensive energy [2, 60]. The interaction is chosen to be ferromagnetic. The $\tau$ variables can be chosen to be spins or rotor variables, but this is not essential for our results.
Many of the fully-connected models studied in the past can be seen as particular instances of Hamiltonian Eq. (1): for example the LMG model [61, 62, 56] is recovered with the choice $H_{0}(\mathbf{s})=-hs_{x}$ and $H_{1}=s_{z}$. Since the interaction is fully invariant under any permutation, we assume the initial state |$\ket{\Psi(0)}$ at $t=0$ to be fully symmetric as well: said |$\ket{k_{j}}$, $j=1,\dots,N$ a collection of single-site states, we can express it as
|$\ket{\Psi(0)}=\frac{1}{\sqrt{N!}}\sum_{P}\bigotimes_{j}\ket{k_{P(j)}}\ ,$ (2)
where the sum runs all over the permutations. In the thermodynamics limit, the mean-field description of the system becomes exact [52, 53, 54, 55, 41, 42]. This means that, as long as we are interested in single-site operators, the dynamics of the system can be described (see [63] for a physical argument) in terms of a single-site effective density operator $\rho(t)$, evolving through the Hamiltonian
$\mathcal{H}(t)=H_{0}(\mathbf{s})-\mu(t)H_{1}(\mathbf{s}),$ (3)
where $\mu(t)$ is determined self consistently via
$\mu(t)=\mathrm{tr}{\left(\rho(t)H_{1}\right)}\ .$ (4)
At $t=0$ we have $\rho(0)=1/N\sum_{j}\ket{k_{j}}\bra{k_{j}}$ |. For finite $N$, the above picture breaks down on timescales polynomial in $N$ (see Ref. [44]). If $\mu(t)=0$, however, the many-body fluctuations dominates over the expectation value of $H_{1}$ regardless of the value of $N$, so that the mean-field approximation is no longer valid.
Let us notice that the dynamics described by Eqs. (3) and (4) conserves the quantity $\epsilon=\mathrm{tr}(\rho(t)H_{0})-\mu^{2}(t)/2$ (see [63]) which corresponds to a physical energy density. As the dynamics is effectively one-body, the long-time behavior of the observable is not expected to converge to the thermal average. In analogy with the classical case, violent relaxation shall occur if $\lim_{t\to\infty}\mu(t)\rightarrow\mu_{\infty}$ which will differ from the thermal expectation value $\mu_{\infty}\neq\mu_{\rm th}$. While violent relaxation is widespread in classical long-range physics, in the quantum realm other scenarios are possible: for example, in the LMG, the periodic oscillations of the magnetization survive at all times in the thermodynamic limit. Even more complicated, non periodic, behaviors are in principle possible. In this Letter we show how, in fact, the presence of violent relaxation in a quantum system can be seen as an exception, rather than as the rule. This is a consequence of the following property, which is the main result of our Letter:
Figure 1: Color plot of the of the magnetization $\mu(t)$ of the two-component quantum rotor model, as a function of the coupling $\lambda$ for a given pure initial state |$\ket{\Psi(0)}\propto 3\ket{L=0}+\ket{L=-1}+\ket{L=1}$. Although the quantitative features of the dynamics of $\mu(t)$ exhibit a sharp transition around $\lambda=0.6$, it is clear from the figure that no violent relaxation actually occur for any value for $\lambda$.
A necessary condition for the system described by mean-field dynamics to undergo violent relaxation is that the (dynamically accessible) point spectrum of the single-site Hamiltonian $\mathcal{H}(t)$ in Eq. (3) contains, at most, a single element (possibly degenerate). In case some of the corresponding states are not accessible to the dynamics because of symmetry reason, those should not be included in the above analysis. If the discrete spectrum is empty, then $\mu_{\infty}=0$.
While a detailed proof of the above statement is provided in the Suppl. Mat. [63], its interpretation is straightforward: in case of violent relaxation $\mathcal{H}(t)$ will become time-independent at large times. Then, the existence of more than one element in the dynamically accessible point spectrum, would result in infinitely-lived Rabi oscillations between the populations of the (two or more) levels in the point spectrum. As a consequence, any system with a finite Hilbert site per site (i.e., in our setting, for any finite $s$) do not display violent relaxation. This surprising outcome underlines the deep difference between classical and the quantum dynamics already for mean-field models.
Examples: In this paragraph we are going to verify our statement by examining some particular instances of the Hamiltonian (1). First, we consider the two-component quantum rotor model [64, 3], whose Hamiltonian can be written as
$H_{\rm QR}=\frac{1}{2}\sum_{j}L^{2}_{j}-\frac{\lambda}{2N}\sum_{j,k}\mathbf{n}_{j}\cdot\mathbf{n}_{k}$ (5)
where $n^{\alpha}_{j}$ ($\alpha=1,2$) are the components of a unity vector ($\mathbf{n}_{j}^{2}=1$) and $L_{j}$ the corresponding angular momenta: $[L_{k},n^{\alpha}_{j}]=i\delta_{jk}\epsilon^{\alpha\beta}n^{\beta}_{k}$ (we set $\hbar=1$). Let us notice that this model can be seen as a large $s$ limit of an Hamiltonian of the form (1). Introducing the angular parametrization $\mathbf{n}_{j}=(\cos\theta_{j},\sin\theta_{j})$, with $[\theta_{j},L_{k}]=i\delta_{jk}$, we can write the single-site Hamiltonian (3) as
$\mathcal{H}_{\rm QR}(t)=\frac{L^{2}}{2}-\lambda\mu_{1}(t)\cos\theta-\lambda\mu_{2}(t)\sin\theta$ (6)
with $\mu_{1}(t)=\mathrm{tr}(\rho(t)\cos\theta)$, $\mu_{2}(t)=\mathrm{tr}(\rho(t)\sin\theta)$. In the following, we will assume a $\theta\rightarrow-\theta$ symmetry in the initial state, so that $\mu_{2}(t)\equiv 0$. The main reason behind the choice of the Hamiltonian (5) is that, in the $\hbar\rightarrow 0$ limit, it reduces to the HMF model [24]. Indeed, while the Hamiltonian Eq. (6) does not exactly coincide with the one of the quantum HMF model [58], we notice that that, for uniform initial conditions, $\rho(t)$ is a pure state and the evolution of the system can be expressed though the Gross-Pitajevskij formalism as in the bosonic HMF model [28, 59].
Figure 2: Color plot of the order parameter $\mu(t)$ of the $w$-model as a function of the coupling $\lambda$ for $w=0$ (left) and $w=1$ (right). The parameter chosen are $h=1$, $s=150$, while the initial state is a pure state with $\braket{m}{\Psi(0)}\propto e^{-m^{2}/4}$. While for $w=0$ $\mu(t)$ converges to a constant asymptotic value regardless of the value of $\lambda$, for $w=1$ at large enough $\lambda$ the system exhibits persistent oscillation, signaling the presence of two bound-state and the absence of violent relaxation. In agreement with our results, the violent-relaxation phases fall within the region $\lambda\mu<x_{c}\sim 1.51$ (left to the white dotted line), in which the single particle Hamiltonian (8) has only one accessible bound states. In both panels, at the left of the violent relaxation phase a delocalized thermal phase occurs where $\lim_{t\to\infty}\mu(t)\to 0$. See the text and the Supp. Mat. [63] for an extended discussion.
In spite of the fact that the single-space Hilbert space is infinite dimensional, let us notice that the spectrum of an Hamiltonian of the form (6) has a completely discrete spectrum for any value of the $\mu$, so that we expect the quantum system not to undergo violent relaxation. This expectation is indeed verified by a numerical analysis of the dynamics of Eq. (6). As shown in Fig. 1, indeed, contrary to the classical case, the evolution of $\mu_{1}(t)$ does not converge to an asymptotic constant value.
In order to see an instance of violent relaxation in the quantum world we introduce the $w$-model Hamiltonian
$H_{w}=\frac{h}{s}\sum_{j}s^{x}_{j}-\frac{\lambda}{2N}\sum_{j,k}\theta_{H}(w^{2}-s^{z2}_{j})\theta_{H}(w^{2}-s^{z2}_{k})$ (7)
which is $\mathbb{Z}_{2}$ symmetric ($s_{j}^{z}\rightarrow-s_{j}^{z}$). In this expression we denote with $\theta_{H}(x)$ the Heaveside step function(with the convention $\theta_{H}(0)=1$) while $w$ is a non-negative integer. Physically, the model is characterized by the competition between the single operator terms and the collective interaction, which tends to localize the system in a subspace with small magnetic quantum number $|m|\leq w$. The corresponding single-site Hamiltonian reads
$\mathcal{H}_{w}(t)=\frac{h}{s}s_{x}-\lambda\mu(t)\theta_{H}(w^{2}-s_{z}^{2}),$ (8)
where $\mu(t)=\mathrm{tr}{\left[\rho(t)\theta_{H}(w^{2}-s_{z}^{2})\right]}$. In the large spin limit $s\to\infty$ and for constant $\mu(t)=\bar{\mu}>0$ the Hamiltonian (8) has a continuum of eigenstates with energy $\mathcal{E}>-h$, delocalized with respect to the magnetic number $m$, and a finite number of non-degenerate bound states, localized around $m=0$, with a discrete set of energies $\mathcal{E}_{n}<-h$. The number $n_{\rm b}$ of the bound states varies from $1$, for $\lambda\bar{\mu}\rightarrow 0^{+}$ to a maximum of $12w+$ 1 in the limit $\lambda\bar{\mu}\rightarrow+\infty$, their parity with respect to $\mathbb{Z}_{2}$ ($m\rightarrow-m$) being $(-1)^{n}$, $n=0,\cdots,n_{\rm b}$ (see [63]). In the following, we will restrict to $\mathbb{Z}_{2}$-even initial conditions, so that the maximum number of accessible bound states is $w+1$.
The $w$ model is the ideal playground to check the link between the discrete spectrum and the presence of violent relaxation. For $w=0$ the discrete spectrum has one element regardless of the value of the parameters. The numerical results shown in Fig. 2 (left panel) (see [63] for the details) demonstrate how the presence of a single bound-state leads to a dynamical phase transition between an asymptotically thermal phase at small values of $\lambda$ and a violent relaxation phase at $\lambda\gtrsim 1.2$. Indeed, at small $\lambda$ $\mu(t)$ vanishes in the long-time limit and the thermal behaviour is recovered. On the contrary, for $\lambda\gtrsim 1.2$ the system undergoes violent relaxation and the system reaches a quasi-stationary state, where $\mu(t)$ attains a non-thermal stationary value $\mu_{\infty}$ at large times. Our analysis does not rule out the presence of small, long-lived oscillations in the localized phase which, anyway, would be compatible with our exact statement.
For $w>1$ a second even bound state emerges for $\lambda\mu>x_{c}(w)$. Its emergence, as shown in Fig. (2) (right panel), results into a new phase for large $\lambda$ in which $\mu(t)$ no longer relaxes; rather it exhibits undamped oscillations. In agreement with our theoretical analysis, the phase in which the system exhibits violent-relaxation falls within the region of parameter space in which $\mathcal{H}_{w}$ has only one accessible bound states, namely $\lambda\mu<x_{c}(1)\sim 1.51$ (see [63]).
Classical limit: The above analysis suggests that violent relaxation should emerge in the classical limit, regardless of the single-particle spectrum of the quantum problem. To understand the mechanism behind this phenomenon let us consider a system with a well-defined classical limit, e.g. the two-component quantum rotor model of Eq. (5). In this case it is convenient to resort to the Wigner function formalism[65, 66, 63]. Let us then introduce the Wigner quasi-probability distribution $W_{t}(x,p)$ relative to the single-particle density operator $\rho(t)$, where $x$ and $p$ are a pair of canonical conjugate variables (e.g. $\theta$ and $L$ in Eq. (6)). The evolution equation
$i\partial_{t}W_{t}(q,p)=\{\{\mathcal{H}_{W}(t),W_{t}(q,p)\}\}\equiv\mathcal{M}_{\mathcal{H}}(W_{t})$ (9)
[…]
By looking at the classical limit of the operator $\mathcal{M}_{\mathcal{H}}$, we are able to show that, under very general assumption on the form of $\mathcal{H}$, that the classical system behaves effectively as a quantum system with a single (infinitely degenerate) element in the discrete spectrum (see [63] for the details). As a consequence, the necessary condition for the violent relaxation is generally met in the classical limit, thus explaining the ubiquitousness of the phenomenon in the classical world.
Conclusions: In this Letter we have explored the possibility for a quantum model with attractive, fully-connected, interactions, to undergo violent-relaxation. While this is the standard behavior in the classical limit, we pointed out that is rather exceptional in the quantum regime due to the presence of discrete points in the spectrum of the effective, mean-field, Hamiltonian. If two or more discrete points are present, resonances between these states give rise to non-decaying oscillations, thus preventing the relaxation of the system. To test such prediction we introduced a new spin model (7), characterized, in the large-$s$ limit, by a tunable number of bound states. We also checked that violent relaxation is absent in the two-component quantum rotor model [64] (5) which, in the classical limit, reduces to the paradigmatic Hamiltonian-mean-field model [24]. Finally, we shown that the necessary conditions in order to have violent-relaxation in the quantum regime, are met under very mild hypothesis in the classical limit.
It is worth noting that Hamiltonian (6) has purely attractive interactions and our findings differ from the ones depicted in Ref. [59], where bi-clustering phenomena have been discussed. However, while our analysis is carried out in the formalism of commuting lattice variables, the fact that the effective Hamiltonian (6) reproduces the Gross-Pitajevskii formalism discussed in Refs. [59, 58] suggests that our picture can be extended to the bosonic case.
All together our results indicate that, regardless of the mean-field nature of such models, quantum effects play a crucial role in the dynamics, and can alter their behaviour already at a qualitative level. As cold atoms into cavity are known to give raise to all-to-all effective interactions, our prediction could be verified experimentally in atomic, molecular, optical setups. On the other hand, the possibility of observing violent-relaxation in quantum model could provide a new way for engineering stationary states in quantum technological applications.
Acknowledgements: N.D. acknowledges funding by the Swiss National Science Foundation (SNSF) under project funding ID: 200021 207537 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster).